Which of the following pH values are within the buffer range for a buffer containing 0.5 M CH3COOH and 0.5 M CH3COONa? The Ka for CH3COO– is 1.8×10–5. Select any answers that apply. Multiple tries are permitted; however, 25% (1/4) point will be deducted for each incorrect response.

Select one or more:

2.76

5.09

0.57

3.81

8.11
Answer #1

Ka = 1.8*10^-5

pKa = - log (Ka)

= - log(1.8*10^-5)

= 4.75

use:

pH = pKa + log {[conjugate base]/[acid]}

= 4.75+ log {0.5/0.5}

= 4.75

Buffer capacity is pH-1 to pH+1

That is from 3.75 to 5.75

pH in this range are:

5.09, 3.81

Answer: 5.09, 3.81